If one card is drawn from a regular desk cards what is the probability the card will not be a ace?
There are total of 52 cards in a deck and 4 of those are Aces. Therefore, there are 48 card non-ace. The probability of drawing one card that is not an ace will be 48 divided by 52. The answer is 12/13.
In maths What is Mutually exclusive probability?
Mutually exclusive means they are independent of one another. So, the two events are independent of one another.
Probability is the what that something will or will not occur?
Probability is the likelihood that something will occur. If you subtract it from 1, we get the likelihood (or probability) that it will not occur.
If a coin is tossed and rolls heads 6 times, the (empirical) probability of obtaining a head is 6/10 or .6.
1-.6 =.4 is the empirical probability (or likelihood) of not getting a head.
Definition of additive law in probability?
This has to do with the union of events. If events A and B are in the set S, then the union of A and B is the set of outcomes in A or B. This means that either event A or event B, or both, can occur.
P(A or B) = P(A) + P(B) - P(A and B)
**P(A and B) is subtracted, since by taking P(A) + P(B), their intersection, P(A and B), has already been included. In other words, if you did not subtract it, you would be including their intersection twice. Draw a Venn Diagram to visualize.
If A and B can only happen separately, i.e., they are independent events and thus P(A and B) = 0, then,
P(A or B) = P(A) + P(B) - P(A and B) = P(A) + P(B) - 0 = P(A) + P(B)
a six
Relate probability to the study of genetics?
Probability is what chance something has to happen. The Punnett Square is a way how to predict in genetics how likely an offspring is to have a trait passed on from parents, or in other words find out the probability of a trait being in the phenotype or the genotype.
What is the probability of drawing a Queen of Clubs or a Queen of Spades from a deck of 52 cards?
Reason:
There are only 2 cards in the deck of 52 which could be either a Queen of Clubs or Queen or Spades, So there is a 2 out of 52 chance, or 3.8461538 % chance.
P(R on the third draw) = 3/5 = 0.60 = 60%
To calculate the probability of this event you have to consider all the possible outcomes that give a Red ball on the third draw. They are four:
[1R,2R,3R], [1R,2W,3R], [1W,2R,3R], [1W,2W,3R].
The probability for these four events are:
1.- P(1R,2R,3R) = (6/10)(5/9)(4/8) = 1/6
2.- P(1R,2W,3R) = (6/10)(4/9)(5/8) = 1/6
3.- P(1W,2R,3R) = (4/10)(6/9)(5/8) = 1/6
4.- P(1W,2W,3R) = (4/10)(3/9)(6/8) = 1/10
Then the probability looked for is the sum of all four:
P(R on the third draw) = 1/6 + 1/6 + 1/6 + 1/10 = 3/5 = 0.60 = 60%
What is the probability of getting five tails in a row?
The probability of getting five tails in a row is 1/2^5, or 1 in 32.
The probability of getting five heads in a row is 1/2^5, or 1 in 32.
Thus, the probability of getting either five heads or five tails in five tosses is 1 in 16.
(The caret symbol means "to the power of," as in 2^5 means "2 to the 5th power.")
0.97
How many face cards in a deck of 52 cards?
There are 4 jacks, 4 queens, and 4 kings or 12 face cards in a deck of 52 cards.
What are some advantages and disadvantages of multistage sampling?
a disadvantige is that a rocket gets disposed of after each flight
Give the example of why probabilities of A given B and B given A are not same?
Let's try this example (best conceived of as a squared 2x2 table with sums to the side). The comma here is an AND logical operator. P(A, B) = 0.1 P(A, non-B) = 0.4 P(non-A, B) = 0.3 P(non-A, non-B) = 0.2 then P(A) and P(B) are obtained by summing on the different sides of the table: P(A) = P(A, B) + P(A, non-B) = 0.1 + 0.4 = 0.5 P(B) = P(A,B) + P(non-A, B) = 0.1 + 0.3 = 0.4 so P(A given B) = P (A, B) / P (B) = 0.1 / 0.4 = 0.25 also written P(A|B) P(B given A) = P (A,B) / P (A) = 0.1 / 0.5 = 0.2 The difference comes from the different negated events added to form the whole P(A) and P(B). If P(A, non-B) = P (B, non-A) then P(A) = P(B) and also P(A|B) = P(B|A).
This is a sampling problem where the entire population is selected and there is no replacement. The chance of selecting all cities in alphabetical order is: (1/10)*(1/9)*(1/8)*(1/7)*(1/6)*(1/5)*(1/4)*(1/3)*(1/2) of 1/(10!) = 2.75 e-7. If we consider alphabetical order can be highest to lowest or lowest to highest, then the probability doubles or: 2/(10!)
What are the characteristics of a normal distribution curve?
Characteristics of a Normal Distribution
1) Continuous Random Variable.
2) Mound or Bell-shaped curve.
3) The normal curve extends indefinitely in both directions, approaching, but never touching, the horizontal axis as it does so.
4) Unimodal
5) Mean = Median = Mode
6) Symmetrical with respect to the mean
That is, 50% of the area (data) under the curve lies to the left of
the mean and 50% of the area (data) under the curve lies
to the right of the mean.
7) (a) 68% of the area (data) under the curve is within one
standard deviation of the mean
(b) 95% of the area (data) under the curve is within two
standard deviations of the mean
(c) 99.7% of the area (data) under the curve is within three
standard deviations of the mean
8) The total area under the normal curve is equal to 1.
How is Probability used in the gambling industry?
It depends on what you mean, but here is a basic description of how it works. Payouts in gambling games are based on the probability of winning, but the payout is less than the true probability of winning. Just as a basic example, the odds of rolling a 12 in the game Craps is 1 in 36, but if bet on the 12 and win, you only win 30 times the amount bet. So basically, you would have to make a $1 bet 36 times to win $30.
Draw five cards without replacement probability of a heart?
As U KNOW THERE ARE 13 CARDS OF HEART IN A DECK AND THERE IS A TOTAL OF 52 CARDS . sO EVERYTIME U TAKE A HEART IT AFFECTS THE TOTAL NO OF CARDS . so THE ANSWER WOULD BE (13*12*11*10*9*8)/(52*51*50*49*48) = 8.42890. app.
I tried my best to explian sorry about the mistakes and i hope it would have helped u
What are the chances of drowning in a bath tub?
The chances of drowning in a bath tub depend on several things, most obvious of which would be your level of fatigue and/or the water level of a drawn bath - If said person is tired and fall asleep, they may eventually slide under the water level, although, in turn, this would wake said person, and through instinct they would rise above the water. Unless someone was fully unconcious, it isn't likely that someone could drown.
What does the Probability of success in a binomial trial symbol?
In typical notation, "p" is the probability of sucess and "q" is the probability of failure. So q = 1 - p. But for your question: p = p.